OpenStudy (alante):

Whats the remainder or(6x3 + 2x2 + 2)divide(x + 2)

OpenStudy (anonymous):

i think its somthing like this (6x3+2x2+2)(d)ivid(2.718282)(x+2) =((6x3+2x2+2)(d)ivid(2.718282))(x+2) =((6x3+2x2+2)(d)ivid(2.718282))(x)+((6x3+2x2+2)(d)ivid(2.718282))(2) =16.309691d2i2vx4+5.436564d2i2vx3+5.436564d2i2vx+32.619382d2i2vx3+10.873127d2i2vx2+10.873127d2i2v =16.309691d2i2vx4+38.055946d2i2vx3+10.873127d2i2vx2+5.436564d2i2vx+10.873127d2i2v

OpenStudy (alante):

@aloud so its 16.3

OpenStudy (anonymous):

no its........ vx+10.873127d2i2v

OpenStudy (alante):

@aloud 10.8

OpenStudy (anonymous):

yup

OpenStudy (alante):

@aloud thats the remainder??? cause i git something different